OpenStudy (anonymous):

find x (x-5)/(3-x)<0 and no the answer isn't x<5

OpenStudy (anonymous):

u say x<5 lets take 4 then we will have (-1)/(-1) and this is equal to +1 after that we can see it is not correct because 1>0

OpenStudy (anonymous):

yes i know that, thats why I said x<5 isn't an answer

OpenStudy (anonymous):

if the answer is x<5 i think the answer is wrong

OpenStudy (anonymous):

but it IS NOT (and i mean NOT) the answer gesh read it right

OpenStudy (anonymous):

i couldnt understand you sorry

OpenStudy (anonymous):

it's x>5

OpenStudy (anonymous):

how did you get that? ( i need to show work + understand)

OpenStudy (anonymous):

times both sides by (3-x) and since you are multiplying by a negative x, the inequality changes the sign. the rest is easy.

OpenStudy (anonymous):

did you get this?

OpenStudy (anonymous):

\[\frac{x-5}{3-x}<0\]

OpenStudy (anonymous):

answer is \[(-\infty,3)\cup(5,\infty)\]

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

consider the two factors separately. \[x-5<0\] if \[x<5\] and \[3-x<0\] if \[x>3\] now compare. if \[x<3\] then \[x-3>0\] and 0\[x-5<0\] and a positive times a negative is negative

OpenStudy (anonymous):

my guess is this is not clear so lets make a chart

OpenStudy (anonymous):

x - 5 ------------------------ 5 +++++++++++ 3-x ++++++++++3 --------------------------- together --------------3+++++++5----------------

OpenStudy (anonymous):

yeah u guess right

OpenStudy (anonymous):

so if x < 3 it is negative, and if x > 5 it is negative, and between 3 and 5 it is positive

OpenStudy (anonymous):

you can also do the following. pick a number less than 3 and test by replacing in the expression. it will be negative pick a number betwen 3 and 5 and test. it will be positive pick a number greater than 5 and test, it will be negative btw i hope it is clear than saying something is < 0 is the same as saying "negative"

OpenStudy (anonymous):

thanks a bizillion